# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v2

@@ 3.2.3 Pasteur’s Quadrant: Revenue Equivalence Theorem and Vickey Auction
In this paper, Vickrey not only expanded fundamental understanding by establishing the earliest version of the Revenue Equivalence Theorem, [^21] Klemperer (1999) provides the following definition of the Revenue Equivalence Theorem: Assuming a fixed number of risk-neutral potential buyers, each with a privately known signal independently drawn from a common, strictly increasing, and atomless distribution, any auction mechanism that satisfies the following conditions—(1) the object is always allocated to the buyer with the highest signal, and (2) a bidder with the lowest feasible signal expects zero surplus—yields the same expected revenue under the Bayesian Nash equilibrium of the resulting game form. one of the foundational results in auction theory, but also introduced the second-price sealed-bid auction, the simplest case of the auction formats now known as the Vickrey auction. [^22] A Vickrey auction is a static auction format for selling any number of identical items, in which the highest bidders each win one item and pay a price equal to the highest losing bid. Interestingly, he devised the second-price sealed-bid auction as a minimalist reform of the first-price sealed-bid auction.
@@ 3.3 Rethinking Innovation Frameworks: Discovery-Invention Cycle
Minimalist market design integrates use-inspired basic research with policy initiatives, fitting well within Pasteur’s Quadrant. However, it aligns even more closely with another framework that challenges Bush’s linear model: the Discovery-Invention Cycle (DIC) model proposed by Narayanamurti and Odumosu (2016). The DIC model shapes the structure of this monograph, organizing the evolution of innovations across diverse applications in Sections 5 through 12 and highlighting the “discovery-invention cycles” connecting them.
@@
My experiences in research and policy since the late 1990s strongly resonate with these ideas. As we progress through the applications in Sections 5 through 12, we will observe how these concepts are reflected in the relationships between various market design innovations, both within specific applications and across them. This will illustrate the strong alignment between minimalist market design and the DIC model.
@@
Drawing inspiration from industrial laboratories like Bell Labs, Narayanamurti and Odumosu argue that Nobel Prize-winning industrial research undermines Bush’s notions of applied and basic research. They highlight that in the industrial research environment, “the very best discovery-oriented research was carried out.” Emphasizing the importance of connections between research and practice, they assert that the most productive distinction was not between basic and applied research but between various research activities and product development.
@@ 4.1 Pursuit of Promising Applications
Minimalist market design is particularly valuable when stakeholders have clearly defined objectives. To assess its potential, consider the following three questions:
@@
    Is there a clear need for a change in the existing system, or is it merely an aspiration of an outsider?
@@
    Does the proposal entail introducing a new system or modifying an existing one?
@@
    Are all key stakeholder goals aligned with neoclassical economics, or do they include essential objectives beyond its scope?
The framework is especially effective when the design economist acts as an external critic intervening in an existing institution, focusing on limitations that extend beyond the textbook considerations of neoclassical economics. Its cautious approach is particularly valuable in areas outside the typical domain of economists, where there is a risk of being seen as intrusive.
@@
4.2 A Guide on the Execution of Minimalist Market Design
@@ 5.2 Housing Markets and the Core
But for me, the most amazing thing about this model is the following procedure—the Gale’s Top Trading Cycles (G–TTC) algorithm—which generates the unique outcome of a housing market. This algorithm was famously suggested to Herbert Scarf by David Gale during one of his early presentations of the housing markets model (Scarf, 2009).
@@ 5.6.2 Contrast with a Technocratic Design Based on the Core
In early 1999, after three years at the University of Michigan in Ann Arbor, I returned to my native Turkey to start as an Assistant Professor at Koç University in İstanbul. Shortly after, in the fall of 2000, we recruited Ünver —a recent Ph.D. graduate from the University of Pittsburgh—for our group. Koç University was relocating to its new campus in Sarıyer, and Ünver would stay in the faculty housing there. During the transition period, he spent a few days at my place.
@@ 6.6 Failed Policy Attempt in Turkey and Lessons Learned
In 1997, even before submitting the paper for scholarly review, I shared my discovery with Turkish officials. Following several correspondences by mail and a meeting with the head of the centralized clearinghouse, ÖSYM,494949ÖSYM is the acronym for Öğrenci Seçme ve Yerleştirme Merkezi, which translates to English as the Student Selection and Placement Center. I received a formal letter from Ankara politely rejecting my proposal. [^50] See Figure 1 for a copy of the letter from the head of ÖSYM rejecting my proposal to replace MCSD with DA in Turkish college admissions, and Figure 2 for its English translation.
@@
Based on this discovery, I decided to explore the formal implications of relaxing the no justified envy axiom. In my view, the primary difference between student placement in Turkish colleges and school choice was the potential dispensability of no justified envy in the latter setting. The flexibility to drop or relax this axiom was valuable to me due to the incompatibility between no justified envy and Pareto efficiency. Naturally, my next objective was to design a Pareto efficient mechanism without sacrificing any other plausible property of DA, such as strategy-proofness or respect for priority improvements. [^51] For school choice, Pareto efficiency already implies individual rationality and non-wastefulness.
@@ 6.11 The 2003 NYC High School Reform: Persuasion Takes a Backseat
Based on these failures, the NYCDOE Director of Strategic Planning reached out to Alvin Roth in May 2003, inquiring whether the matching process used for the U.S. medical match could be adapted for designing a new high school matching process. Like the DA mechanism, the Roth and Peranson design (see Section 3.2.2), used for the U.S. medical match, are also based on Gale and Shapley’s celebrated individual-proposing Deferred Acceptance algorithm.
@@ 7 Kidney Exchange
In the mid-2000s, Alvin Roth, Utku Ünver, and I laid the theoretical foundation for organized kidney exchange through a series of papers (Roth et al., 2004, 2005b, 2007). Leveraging formal techniques from market design and optimization, kidney exchange rapidly revolutionized living donor kidney donation in numerous countries, saving thousands of lives annually within just a few years (Purtill, 2018; Rose, 2019). As an unexpected yet transformative application of market design, these developments played a pivotal role in elevating the field’s practical relevance and visibility during this period.
The recognition for these efforts reached a milestone in 2012 when our co-author, Alvin Roth, was honored with the Nobel Memorial Prize in Economic Sciences, jointly with Lloyd Shapley. Notably, kidney exchange was prominently featured in Torsten Persson’s award ceremony speech as a committee member (Persson, 2012).
But how did a team of economists manage to develop the tools and establish the infrastructure that now consistently touches so many lives? This section explores the pivotal role of key ideas in minimalist market design, illustrating their substantial influence on the success and widespread adoption of organized kidney exchange clearinghouses worldwide. The policy and practical impacts of kidney exchange and school choice, in turn, later led to my systematic application of these ideas in subsequent market design efforts and, ultimately, to the development of minimalist market design, as formulated in this monograph.
@@ 7.3 First Kidney Exchange Model in Roth, Sönmez, and Ünver (2004)
For a use-inspired theory project, our approach to kidney exchange differed significantly from my approach to school choice. As discussed in Section 6, in school choice, we advocated for a few highly refined policy recommendations. In contrast, at this stage, our focus in kidney exchange was primarily on providing insights into the virtues of:
@@ 7.8 Global Kidney Exchange
I refrain from taking a position on whether the sharp criticisms of GKE by the EU NCA and other international bodies are justified, as many of these objections fall beyond my expertise. What is relevant for my purposes is that GKE faced substantial criticism from various organizations, indicating a level of ineffectiveness in achieving the intended welfare gains from kidney exchange. The contrast between the early success in kidney exchange and the sharp opposition to GKE underscores the importance of adopting a cautious approach—a key feature of minimalist market design—when establishing partnerships with experts in other disciplines and policymakers.
@@ 7.9 Kidney Exchange Worldwide: Two Decades On
This challenge was finally overcome in March 2007, when the Justice Department issued a legal memo stating that kidney exchanges do not constitute “valuable consideration” under the provisions of NOTA, thereby not violating the act. Later that year, in December 2007, the U.S. Senate passed an amendment to NOTA—the Charlie W. Norwood Living Organ Donation Act—explicitly clarifying that kidney exchange is legal in the U.S. With this breakthrough, the pathway was opened for national kidney exchange. [^85] Before the enactment of the Charlie W. Norwood Living Organ Donation Act in the U.S., kidney exchange had already been deemed legal in the UK under the provisions of the Human Tissue Acts of 2004 (covering England, Northern Ireland, and Wales) and 2006 (for Scotland) (Biro et al., 2019). Since then, with active support from members of the market design community in many instances, similar laws have been enacted in several other countries, enabling various forms of kidney exchange. Examples include the Belgian Living Donor Exchange Protocol in Belgium in 2008 (Biro et al., 2019), the Bioethics Law in France—initially in 2011 (allowing only 2-way exchanges) and later expanded in 2021 (permitting exchanges of up to six ways) (Combe et al., 2022), and an amendment to the Transplantation Law in Germany in 2024 (Ashlagi et al., 2024; Ockenfels et al., 2024).
@@ 7.9.2 The Global Rise of Kidney Exchange Programs Since 2008
After the first kidney exchange performed in the U.S. in Rhode Island in 2000, progress in kidney exchange transplantation was initially slow. Between 2000 and 2003—before our engagement with the transplantation community—only 31 kidney exchange transplants were performed nationwide over a four-year span. In stark contrast, 4,387 transplants were conducted during the four-year period from 2020 to 2023 (OPTN, 2024a). (For annual numbers of kidney exchange transplants in the U.S. between 2000 and 2023, see Figure 7.)
In 2023 alone, a record 1,283 kidney exchange transplants were performed in the U.S. By the first week of December 2024, nearly 13,000 patients in the country had received transplants through kidney exchange.
@@
Over the past two decades, organized kidney exchange programs have expanded beyond the U.S., providing life-saving transplants to patients worldwide. Countries such as Turkey, the U.K., Canada, India, Saudi Arabia, the Netherlands, and Australia now perform some of the highest kidney exchange volumes globally outside of the U.S.
More than 80% of kidney transplants in Turkey are from living donors. In 2023, the country ranked first worldwide in the living kidney donation rate, with more than 35 donors per million people (IRODaT, 2024). Turkey also ranks third globally in living-donor kidney transplants (LDKTs), following India and the U.S. Since 2018, over 3,000 transplants have been performed annually, except in 2020, when the number dropped to 2,250 due to the COVID-19 pandemic.
Kidney exchange is widely practiced in Turkey across numerous single-hospital programs, accounting for more than 2,000 transplants as of 2023. However, NDD chains are not practiced due to the prohibition of altruistic donation. The Turkish Ministry of Health and the Turkish Society of Nephrology have tracked kidney exchange activity nationwide since 2012 (Turkish Society of Nephrology, 2023). According to their registries, 1,913 kidney exchange transplants were performed between 2012 and 2022, accounting for 6.5% of the country’s total LDKTs. This relatively modest rate reflects the absence of a national kidney exchange system (see Section 12.4) and the limited availability of such programs in only some hospitals. However, in hospitals participating in kidney exchange, rates appear significantly higher, around 15%. For example, Altun and Yavuz (2023) documented 141 kidney exchange transplants at Istanbul Göztepe Medicalpark Hospital from July 2011 to June 2021, representing 16.1% of the LDKTs performed at the center during that time.
Kidney exchange activity in Turkey predates 2012, with significant progress reported between 2006 and 2011. As documented by Yücetin et al. (2013), among the 59 hospitals performing LDKTs nationwide, Antalya Medicalpark Hospital, Istanbul Göztepe Medicalpark Hospital, and Akdeniz University Medical School Transplantation Center recorded the highest kidney exchange volumes during this period, conducting 272 transplants, which accounted for 14.4% of the LDKTs at these centers.
@@
The U.K. was the first country outside the U.S. to adopt market design tools for kidney exchange. Since 2007, the National Health Service Blood and Transplant (NHSBT) has coordinated kidney exchanges nationally through the National Living Donor Kidney Sharing Scheme. This program, which operates every three months and has collaborated with market design researcher David Manlove since 2008, became the first national system to implement optimal matching procedures. Initially limited to 2-way and 3-way cycles (Biró et al., 2009), the scheme began incorporating NDD chains of up to 3-way in 2012 (Manlove and O’Malley, 2014; Ferrari et al., 2014). The program—the largest of its kind in Europe—celebrated its 1,000th kidney exchange transplant in March 2019.
@@
As the most populous country in the world, India leads globally in LDKTs, with more than 10,000 performed in 2022 (IRODaT, 2024). Kidney exchange is widespread and conducted through several single-hospital programs across the country. Although NDD chains are not practiced in India, as in Turkey, due to the illegality of altruistic donation, kidney exchange operations are often carried out over several days, enabling the execution of large kidney exchange cycles.
Kute et al. (2024) reports that the Dr. HL Trivedi Institute of Transplantation Sciences in Ahmedabad—described by the authors as the largest single-center kidney exchange program in Asia—performed 539 kidney exchange transplants between January 2000 and March 2024, including one 10-way and two 6-way exchanges conducted over multiple days. These transplants accounted for 10% of the 5,346 LDKTs performed at the center during that period. Similarly, Mahatma Gandhi Medical College in Jaipur, identified by Mangal et al. (2024) as the largest program in Northern India, reported conducting 127 kidney exchange transplants between 2014 and March 2024.
In Saudi Arabia, as in Turkey and many other countries in Asia, more than 80% of kidney transplants are from living donors. In 2023, the country ranked second worldwide—after Turkey—in the living kidney donation rate, with more than 35 donors per million people (IRODaT, 2024). The King Faisal Specialist Hospital and Research Centre reported performing its 500th kidney exchange transplant in October 2024 as part of its single-hospital Kidney Paired Donation Program, which has been in operation since 2011.
The Netherlands is the only European country where transplants from living donors are as prevalent as those from deceased donors, with 400–500 transplants annually from each source since the early 2010s. In 2023, the country ranked fourth worldwide—after Turkey, Saudi Arabia, and Israel—in the living kidney donation rate, with more than 28 donors per million people (IRODaT, 2024). With the participation of eight transplant centers and one histocompatibility center, the Netherlands launched its national kidney exchange program in 2004, becoming the first country in Europe to establish such a system (Klerk et al., 2005). The program allows up to 4-way exchanges and facilitated 284 kidney exchange transplants between 2004 and 2016 (Biro et al., 2019). Although altruistic donation is permitted in the country, and NDD chains are performed regionally, these are not included in the national system. Ferrari et al. (2014) reports that 82 transplants were facilitated through regional NDD chains as of April 2014.
@@
In Europe, similar collaborations have emerged to increase transplant opportunities, particularly in countries with relatively small numbers of living donors. One of the most active initiatives is the Scandiatransplant Kidney Exchange Programme (STEP), which involves transplant centers in Denmark, Finland, Norway, and Sweden (Weinreich et al., 2023). Launched in 2019 with the advocacy of design economist Tommy Andersson, STEP initially supported only 2-way exchanges. However, as of January 2024, it now accommodates up to 3-way exchanges and non-directed donor (NDD) chains. Between its inception in 2019 and September 2024, STEP enabled a total of 82 transplants (Scandiatransplant, 2024).

### v3

@@ 3.2.3 Pasteur’s Quadrant: Vickey Auction and Revenue Equivalence Theorem
In this paper, Vickrey not only advanced fundamental understanding by establishing the earliest version of the Revenue Equivalence Theorem, [^22] Klemperer (1999) provides the following definition of the Revenue Equivalence Theorem: Assuming a fixed number of risk-neutral potential buyers, each with a privately known signal independently drawn from a common, strictly increasing, and atomless distribution, any auction mechanism that satisfies the following conditions—(1) the object is always allocated to the buyer with the highest signal, and (2) a bidder with the lowest feasible signal expects zero surplus—yields the same expected revenue under the Bayesian Nash equilibrium of the resulting game form. one of the foundational results in auction theory, but also introduced the second-price sealed-bid auction, the simplest case of what are now known as Vickrey auctions. [^23] A Vickrey auction is a static auction format for selling any number of identical items, in which the highest bidders each win one item and pay a price equal to the highest losing bid. Interestingly, he devised the second-price sealed-bid auction as a minimalist reform of the first-price sealed-bid auction.
@@ 3.3 Rethinking Innovation Frameworks: Discovery-Invention Cycle
Minimalist market design integrates use-inspired basic research with policy initiatives, positioning it firmly within Pasteur’s Quadrant. Yet, it aligns even more closely with another framework that also challenges Bush’s linear model: the Discovery-Invention Cycle (DIC) model by Narayanamurti and Odumosu (2016). This model shapes the structure of this monograph, tracing the evolution of innovations across diverse applications in Sections 5 through 12 and highlighting the “discovery-invention cycles” that connect them.
@@
According to this distinction, much of the work in commissioned market design projects, performed as a scheduled activity leading to the adopted design and “often required to be fast,” according to Roth (2002), is considered “development” rather than “research,” unlike the work performed after its completion as an unscheduled activity. However, in the spirit of Peter Medawar’s concerns discussed in Section 1.2, this research reflects post hoc research progress after the practical or policy impact has already occurred. Thus, it is less clear how the adapted design in commissioned settings may be received in cases where the design economist is an outsider advocate, since it is not stress-tested in such situations. Indeed, this has been a key reason why I sought an alternative approach to market design as a junior economist.
@@
My experiences in research and policy since the late 1990s strongly resonate with these ideas. As we progress through the applications in Sections 5 through 12, we will observe how these concepts are reflected in the relationships between various market design innovations, both within specific applications and across them. This will illustrate the strong alignment between minimalist market design and the DIC model of innovation.
@@
Drawing inspiration from industrial laboratories like Bell Labs, Narayanamurti and Odumosu argue that Nobel Prize-winning industrial research undermines Bush’s notions of applied and basic research. They highlight that in the industrial research environment, “the very best discovery-oriented research was carried out.” Emphasizing the importance of connections between research and practice, they assert that the most productive distinction was not between basic and applied research but between various unscheduled research activities and scheduled product development.
@@ 4.1 Pursuit of Promising Applications
Minimalist market design proves most valuable when stakeholders have well-defined objectives, and the existing system, while imperfect, is not so fundamentally broken that it cannot be substantially preserved. [^29] An example of a system too difficult to salvage with a minimalist approach is the queuing system used by the non-profit Feeding America before 2005 to allocate excess food across food banks. This system neither disclosed the type of food being allocated nor accounted for the specific needs of individual food banks, resulting in severe welfare loss. Recognizing these significant shortcomings, and with guidance from four University of Chicago faculty members, the organization replaced the queuing system with an elaborate auction run twice daily. In this new system, each food bank made “purchases” using a specialized currency called “shares,” which were determined for each auction based on the bank’s characteristics and purchase history (Prendergast, 2017, 2022). To assess its potential, consider the following three questions:
@@
    Are you commissioned by decision makers or other stakeholders seeking changes to the existing system, or is it merely an aspiration of an outsider?
@@
    Does the proposal involve introducing a new system or modifying an existing one?
@@
    Are all key stakeholder goals within the purview of economics, or do they include essential objectives beyond its scope?
This framework is especially useful when the design economist acts as an external critic, intervening in an existing institution that, while flawed, does not require a complete overhaul, and addressing its limitations that go beyond the textbook considerations of economics. Its cautious approach is particularly valuable in areas outside the traditional domain of economists, where interventions risk being perceived as intrusive.
@@
4.2 A Roadmap for an Aspiring Market Designer
@@ 5.2 Housing Markets and the Core
But for me, the most amazing thing about this model is the following procedure—the Gale’s Top Trading Cycles (G–TTC) algorithm—which generates the unique outcome of a housing market. This algorithm was famously invented by David Gale and suggested to Herbert Scarf during one of Scarf’s early presentations of the housing markets model (Scarf, 2009).
@@ 5.6.2 Contrast with a Technocratic Design Based on the Core
In early 1999, after three and a half years at the University of Michigan in Ann Arbor, I returned to my native Turkey to start as an Assistant Professor at Koç University in İstanbul. Shortly after, in the fall of 2000, we recruited Ünver —a recent Ph.D. from the University of Pittsburgh—for our group. Koç University was relocating to its new campus in Sarıyer, and Ünver would stay in the faculty housing there. During the transition period, he spent a few days at my place.
@@ 6.6 Failed Policy Attempt in Turkey and Lessons Learned
In 1997, even before submitting the paper for scholarly review, I shared my discovery with Turkish officials. Following several correspondences by mail and a meeting with the head of the centralized clearinghouse, ÖSYM,515151ÖSYM is the acronym for Öğrenci Seçme ve Yerleştirme Merkezi, which translates to English as the Student Selection and Placement Center. I received a formal letter from Ankara politely rejecting my proposal. [^52] See Figure 1 for a copy of the letter from the head of ÖSYM rejecting my proposal to replace MCSD with DA in Turkish college admissions, and Figure 2 for its English translation.
@@
Based on these findings, I decided to explore the formal implications of relaxing the no justified envy axiom. In my view, the primary difference between student placement in Turkish colleges and school choice was the potential dispensability of no justified envy in the latter setting. The flexibility to drop or relax this axiom was valuable to me due to the incompatibility between no justified envy and Pareto efficiency. Naturally, my next objective was to design a Pareto efficient mechanism without sacrificing any other plausible property of DA, such as strategy-proofness or respect for priority improvements. [^53] For school choice, Pareto efficiency already implies individual rationality and non-wastefulness.
@@ 6.11 The 2003 NYC High School Reform: Persuasion Takes a Backseat
Based on these failures, the NYCDOE Director of Strategic Planning reached out to Alvin Roth in May 2003 to explore whether the matching process used for the U.S. medical residency match could be adapted to design a new high school matching process. Similar to the DA mechanism advocated by Balinski and Sönmez (1999) for student placement and by Abdulkadiroğlu and Sönmez (2003) for school choice, the Roth and Peranson design (see Section 3.2.2), used for the U.S. medical residency match, is also based on Gale and Shapley’s celebrated individual-proposing Deferred Acceptance algorithm.
@@ 7 Kidney Exchange
In the mid-2000s, through a series of papers, a team of design economists consisting of Alvin Roth, Utku Ünver, and me laid the theoretical foundation for organized kidney exchange (Roth et al., 2004, 2005b, 2007). Leveraging formal techniques from market design and optimization, kidney exchange rapidly transformed living donor kidney donation in numerous countries, saving thousands of lives annually within just a few years (Purtill, 2018; Rose, 2019). As an unexpected yet transformative application of market design, these developments played a pivotal role in elevating the field’s practical relevance and visibility during this period.
Recognition for these efforts reached a milestone in 2012 when our co-author, Alvin Roth, was honored with the Nobel Memorial Prize in Economic Sciences, jointly with Lloyd Shapley. Notably, kidney exchange was prominently featured in Torsten Persson’s award ceremony speech as a committee member (Persson, 2012).
But how did a team of economists manage to develop the tools and establish the infrastructure that now consistently touches so many lives? This section explores the pivotal role of key ideas in minimalist market design, illustrating their profound influence on the success and widespread adoption of organized kidney exchange clearinghouses worldwide. The policy and practical impacts of kidney exchange and school choice, in turn, later inspired my systematic application of these ideas in subsequent market design efforts and ultimately led to the conceptualization of minimalist market design, as presented in this monograph.
@@ 7.3 First Kidney Exchange Model in Roth, Sönmez, and Ünver (2004)
As a use-inspired theory project, our approach to kidney exchange differed markedly from my earlier approach to school choice. In school choice, we offered a set of refined policy recommendations (see Section 6.10). In contrast, at this stage, our focus in kidney exchange was primarily on providing insights into the virtues of:
@@ 7.8 Global Kidney Exchange
I refrain from taking a position on whether the sharp criticisms of GKE by the EU NCA and other international bodies are justified, as many of these objections fall outside my expertise. Indeed, GKE has also received widespread endorsements, including from prominent bioethicists such as Peter Singer (Minerva et al., 2019) and the Italian government, which recommended GKE to the World Health Organization (Pullen, 2018). What is relevant for my purposes is that GKE faced substantial criticism from various organizations, reducing its effectiveness in achieving the intended welfare gains from kidney exchange. The contrast between the early success of kidney exchange and the sharp opposition to GKE underscores the importance of adopting a cautious approach—a key feature of minimalist market design—when establishing partnerships with experts in other disciplines and policymakers.
@@ 7.9 Kidney Exchange Worldwide: Two Decades On
This challenge was finally overcome in March 2007, when the Justice Department issued a legal memo stating that kidney exchanges do not constitute “valuable consideration” under the provisions of NOTA, thereby not violating the act. Later that year, in December 2007, the U.S. Senate passed an amendment to NOTA—the Charlie W. Norwood Living Organ Donation Act—explicitly clarifying that kidney exchange is legal in the U.S. With this breakthrough, the pathway was opened for national kidney exchange. [^87] Before the enactment of the Charlie W. Norwood Living Organ Donation Act in the U.S., kidney exchange had already been deemed legal in the UK under the provisions of the Human Tissue Acts of 2004 (covering England, Northern Ireland, and Wales) and 2006 (for Scotland) (Biro et al., 2019). Since then, with active support from members of the market design community in many instances, similar laws have been enacted in several other countries, enabling various forms of kidney exchange. Examples include the Belgian Living Donor Exchange Protocol in Belgium in 2008 (Biro et al., 2019), and the Bioethics Law in France—initially in 2011 (allowing only two-way exchanges) and later expanded in 2021 (permitting exchanges and NDD chains of up to three-way) (Combe et al., 2022; Biró et al., 2024). In addition, as of December 2024, a draft law for an amendment to the Transplantation Law in Germany allowing for kidney exchange between incompatible pairs as well as NDD chains is pending approval in Parliament (Ashlagi et al., 2024; Ockenfels et al., 2024).
@@ 7.9.2 The Global Rise of Kidney Exchange Programs Since 2008
After the first kidney exchange performed in the U.S. in Rhode Island in 2000, progress in kidney exchange was initially slow. Between 2000 and 2003—prior to our engagement with the transplantation community—only 31 kidney exchange transplants were performed nationwide over four years. In sharp contrast, 4,387 transplants were conducted between 2020 and 2023 (OPTN, 2024a). A record 1,283 kidney exchange transplants were performed in 2023 alone, and by December 2024, the cumulative total exceeded 13,000 patients in the U.S. (For annual transplant numbers from 2000 to 2023, see Figure 7.)
@@
Over the past two decades, kidney exchange programs have expanded beyond the U.S., providing life-saving transplants to patients worldwide. Countries such as the U.K., Turkey, Canada, India, Saudi Arabia, the Netherlands, and Australia now perform some of the highest kidney exchange volumes outside the U.S. Altogether, the number of kidney exchange transplants carried out globally to date exceeds 20,000.
@@
The U.K. was the first country outside the U.S. to adopt market design tools for kidney exchange, including the implementation of optimal matching algorithms, and it now operates the largest national system. Since 2007, the National Health Service Blood and Transplant (NHSBT) has coordinated kidney exchanges through the UK Living Donor Kidney Sharing Scheme. This program, which runs every three months, has collaborated with market design researcher David Manlove since 2008 and celebrated its 2000th kidney exchange transplant in September 2024.
Initially limited to 2-way and 3-way cycles (Biró et al., 2009), the U.K. scheme has also been organizing non-directed donor (NDD) chains of up to three recipients since 2012 (Manlove and O’Malley, 2014; Ferrari et al., 2014). Among the 670 kidney exchange transplants conducted between April 2020 and March 2024, 126 (18.8%) were carried out in 2-way cycles, 118 (17.6%) in 3-way cycles, 215 (32.1%) in 2-way NDD chains, and 211 (31.5%) in 3-way NDD chains (NHS Blood and Transplant, 2024).
Two key factors behind the U.K.’s success are the substantial participation of altruistic non-directed donors (NDDs) and compatible donor-recipient pairs. Between April 2015 and March 2024, there were 653 NDDs, with 347 donating directly to the deceased donor list and 306 donating through the kidney sharing scheme (NHS Blood and Transplant, 2024). Over this period, the proportion of NDDs participating in kidney exchanges grew from approximately 30% to 35% to 55% to 65%. Additionally, 119 compatible pairs participated in kidney exchanges during the same period.
Taken together, the U.K.’s centrally managed national system—which incorporates up to 3-way cycles and NDD chains, employs an optimal matching algorithm, and benefits from significant participation by altruistic donors and compatible pairs—stands out as perhaps the closest large-scale realization of the vision pioneered in Roth et al. (2004, 2005b, 2006, 2007).
In Turkey, over 80% of kidney transplants are from living donors, and the country led the world in living kidney donation rates in 2023, exceeding 35 donors per million people (IRODaT, 2024). Performing over 3,000 transplants annually, Turkey also ranks third globally in living-donor kidney transplants (LDKTs), behind only India and the U.S. These factors make Turkey an ideal candidate to implement a national centralized system. Unfortunately, as discussed later in Section 12.4, my efforts with Ünver in this direction have not succeeded so far.
Nevertheless, kidney exchange is widely practiced in Turkey through several single-hospital programs, with more than 2,000 transplants performed as of 2023. However, NDD chains are not utilized, as altruistic donation is prohibited by law.
The Turkish Ministry of Health and the Turkish Society of Nephrology have tracked kidney exchange activity nationwide since 2012 (Turkish Society of Nephrology, 2023). According to their registries, 2,067 kidney exchange transplants were performed between 2012 and 2023, accounting for 6.3% of the country’s total LDKTs. This relatively modest rate reflects the absence of a national kidney exchange system and the limited availability of kidney exchange in only some hospitals. However, in hospitals with kidney exchange programs, rates are significantly higher, around 15%. For example, Altun and Yavuz (2023) documented 141 kidney exchange transplants at Istanbul Göztepe Medicalpark Hospital from July 2011 to June 2021, representing 16.1% of the LDKTs performed at the center during that period.
Kidney exchange activity in Turkey predates 2012, with significant activity reported between 2006 and 2011. According to Yücetin et al. (2013), among the 59 hospitals performing LDKTs nationwide, Antalya Medicalpark Hospital, Istanbul Göztepe Medicalpark Hospital, and Akdeniz University Medical School Transplantation Center reported the highest kidney exchange volumes during this period. Together, they conducted 272 transplants, accounting for 14.4% of the LDKTs carried out at these centers.
@@
India, the most populous country in the world, leads globally in LDKTs, performing over 10,000 transplants in 2022 (IRODaT, 2024). Kidney exchange is widespread and conducted through numerous single-hospital programs across the country. However, as in Turkey, NDD chains are not practiced because altruistic donation is prohibited by law. Unlike in many other countries, kidney exchange operations in India are often carried out over several days, enabling the execution of large exchange cycles.
Kute et al. (2024) reports that the Dr. HL Trivedi Institute of Transplantation Sciences in Ahmedabad—described as the largest single-center kidney exchange program in Asia—performed 539 kidney exchange transplants between January 2000 and March 2024. These exchanges included one 10-way and two 6-way exchanges conducted over multiple days and accounted for 10% of the 5,346 LDKTs performed at the center during this time. Similarly, Mahatma Gandhi Medical College in Jaipur, described by Mangal et al. (2024) as the largest program in Northern India, reported conducting 127 kidney exchange transplants between 2014 and March 2024.
In Saudi Arabia, as in Turkey and many other Asian countries, more than 80% of kidney transplants are from living donors. In 2023, the country ranked second worldwide—after Turkey—in the living kidney donation rate, with more than 35 donors per million people (IRODaT, 2024). The King Faisal Specialist Hospital and Research Centre celebrated its 500th kidney exchange transplant in October 2024 as part of its single-hospital Kidney Paired Donation Program, which has been in operation since 2011.
The Netherlands is the only European country where transplants from living donors are as prevalent as those from deceased donors, with 400–500 transplants annually from each source since the early 2010s. In 2023, the country ranked fourth worldwide—after Turkey, Saudi Arabia, and Israel—in the living kidney donation rate, with more than 28 donors per million people (IRODaT, 2024).
With the participation of eight transplant centers and one histocompatibility center, the Netherlands launched its national kidney exchange program in 2004, becoming the first country in Europe to establish such a system (Klerk et al., 2005). The program allows up to 4-way exchanges and facilitated 284 kidney exchange transplants between 2004 and 2016 (Biro et al., 2019). Although altruistic donation is legal in the country and NDD chains are performed regionally, they are not managed within the national system. Ferrari et al. (2014) reports that 82 transplants had been facilitated through regional NDD chains as of April 2014.
@@
In Europe, similar cross-border collaborations have emerged over the past decade, particularly in countries with relatively few living donors. One of the most active initiatives is the Scandiatransplant Kidney Exchange Programme (STEP), which involves transplant centers in Denmark, Finland, Norway, and Sweden (Weinreich et al., 2023).
Launched in 2019 with the advocacy of design economist Tommy Andersson and initially limited to 2-way exchanges, STEP has been organizing up to 3-way exchanges and NDD chains since January 2024. As of September 2024, the program has facilitated a total of 82 transplants (Scandiatransplant, 2024).